# Multi-MeV electrons observed by CRRES during solar cycle 22

### v1

@@ 4 HEEF Dataset Analysis
To examine electron lifetimes throughout the CRRES era, Figure 5 shows data in the lowest energy channel, 0.65 MeV, now ordered as a function of radial distance. The fits are once again overlaid on top of the data. The automated algorithm returns increasing numbers of fits at the intermediate L shells examined, overall visibly well-capturing exponentially decaying fluxes. A few flux decreases can however also be seen to be missed. In some instances, e.g. Figure 5c, it might appear that there are two stages to the decay process, for example in Figure 6h after day 450. These often appear as a shallow gradient followed by a greater decrease in the differential number flux. Further non-uniform decays are evident near the peak fluxes themselves over shorter time periods, with notably rounded structures at the lower L shells. This may in part be due to increasing production close to the initial event, varying geomagnetic activity levels during the decay, or potentially different eigenmodes dominating the pitch angle diffusion. At the lowest L shells, the electron fluxes decay over extended periods of tens-hundreds of days. The automated nature of the algorithm struggles with this extended period due to a short minimum fit window required to be defined to resolve the shorter timescales at larger L shells. While tweaks to the fitting procees improve these, this is difficult to scale across all L shells due to the rapid gradient in lifetimes between the inner belt and slot region and assumptions behind exponentially decaying fluxes following an enhancement. We also wish to maintain the automated neutrality of the algorithm and, despite the difficulties in the inner zone, a few fits are returned that are representative of this region. The decay periods are well-represented by the exponential fits in the region 3.0 $\leq$ L* $\leq$ 4.5. At L* $\geq$ 5, only a few fits are visible. This may again in part be due to the algorithm being difficult to optimise for such a wide spread in lifetimes where, at these greater radial distances, further transport processes associated with radial diffusion [Su et al., 2010], drift shell splitting [Sibeck et al., 1987] and drift orbit bifurcations [Desai et al., 2021a] become significant.
Figure 6 displays the results across all L shells and energies with each fit colored according to its mean Kp index. The average fit at each 0.1 L bin is also shown as a black line together with error bars representing the standard deviation.
At the lowest L shells in the inner belt, the lifetimes of the 0.65 MeV electrons are tens to over a hundred days. The difficulty in fitting to these L shells, see Figure 5a and associated discussion, means the fits appear to be biased towards shorter time-periods where the two fits appear of a greater gradient than the seemingly constant decay between day 220 and 325. This appears to have been caused by storms at or prior to the start of the mission followed by relatively quiet conditions for 100 days which resulted in an even longer decay period than those shown, of $\approx$200 days. Given these extended timescales, these fluxes may have partially resulted from the extreme storm of March 1989 [Boteler, 2019]. These long timescales are also evident in that during the second half of the mission, where geomagnetic activity was higher, the 0.65 MeV flux is unable to decay and appears to be largely on an upward trend from around day 450 to  550. The decreased slot region lifetimes between L* = 2-3 are well captured at this energy, as in Figure 3, before the lifetimes increase to another smaller maximum with increasing radial distance near L = 4.
@@
The lifetimes of the 0.95 MeV electrons are lower than those of the 0.65 MeV electrons near L* = 4.0 but this trend does not extend into the slot region where the reverse is true. There is a peak near L = 2.5-3 but this is due to a large spread in the data here. At energies greater than 1 MeV, the peak lifetime in the outer belt moves inward with increasing energy from L* $\approx$ 4.5 at 1 MeV down to L* $\approx$ 3 at 4.55 MeV. The location of the shortest lifetimes in the slot region fluxes also move inwards, with increasing energies displaying longer lifetimes.
@@
In addition to the average trends, the Kp values of each fitted lifetime provides information on how the decays scale with geomagnetic activity. At higher L shells, i.e. greater than L* = 3, elevated geomagnetic activity appears to drive fast losses with some of the shortest timescales showing Kp values near 5. At the lower L’s, however, and particularly at te lowest energy channel of 0.65 MeV, high Kp values often correspond to longer decay timescales. High Kp scenarios will still result in injections from the tail and acceleration of energetic particles and this trend is therefore explained by sustained production of electrons occurring alongside pitch angle scattering, a phenomenon which preferentially affects lower energies. This is further discussed in Section 5.
@@
Figure 7: Experimental and theoretical electron lifetimes as a function of L shell for (a-h) 0.65, 0.95, 1.60, 2.00, 2.35, 2.75, 3.15 and 3.75 MeV electrons respectively. The black, red, grey, green and pink traces represent the experimental lifetimes derived from the CRRES-HEEF (this study), MEA [Meredith et al., 2006], VAP [Claudepierre et al., 2020], SAC/DEMETER [Benck et al., 2010] and SAMPEX datasets [Meredith et al., 2009], respectively. The dashed blue and orange traces represent the pitch angle diffusion theoretical lifetimes for quiet and high geomagnetic activity, respectively. All data is interpolated to the HEEF energies, whereas the SAMPEX 2–6 MeV data are associated with the 2.35 MeV HEEF channel.
@@ 5 Comparisons & Discussion
In Figure 6, the pitch angle diffusion theory matches the observed trend of increased lifetimes near the inner belt at 0.65 MeV and depressed lifetimes in the slot region. In the inner part of the outer radiation belt, the quiet-time theoretical lifetimes at this energy are a factor of two or more smaller than the observed lifetimes in the region 3.2 $<L*<$ 4.6. These larger HEEF lifetimes are similar to those observed by SAC/DEMETER in 2001 [Benck et al., 2010] and also by ATS-1 [Vampola, 1971] see Claudepierre et al. [2020, Fig. 2f therein]. This however is not observed in the VAP lifetimes. Further examination of Figure 6 reveals a cause of this discrepancy. The underlying HEEF fits, most visibly between L* = 3-4, feature a clustering of lifetimes at the lower lifetimes ranges, similar to those reported by the MEA and closer to those observed by VAP. However, an additional clustering above this demonstrates a bi-modality within the dataset, with this second cluster corresponding to higher Kp values. This bimodality is also identifiable in the underlying fits on Figure 5d and 5e where many of the fitted lifetimes have shallower gradients.
@@
The consequent differences in the average lifetimes in Figure 7a may therefore have been caused by differing levels of geomagnetic activity, with CRRES having observed an average Kp index of 2.67 compared to the VAP average of 1.67. This can still be explained through three distinct mechanisms. Firstly, the intensity of geomagnetic storms were significantly greater in solar cycle 22 compared with solar cycle 24. These, during the latter half of the mission, indeed appear to be associated with the slower decays in Figure 5. These intense and longer lasting events also influence the average Kp indices of a decay and thus skew the decay rates to higher Kp values, producing a “left-right" Kp bias as indeed shown by Meredith et al. [2006, Figure 6 therein]. Secondly, disturbed geomagnetic conditions would more frequently reoccur during a decay period. This is visibly consistent with several of the decays observed, see for example the decays days 270 and 295 in Figure 5e-f where large injections/enhancements occur during the decay process. Thirdly, it is well-known that typical solar wind parameters have declined significantly over the past 50 years, with differences between the 70-90s and 2010s having been highlighted as high as tens to over fifty percent [McComas et al., 2013]. Such long-term trends in the radiation belts are difficult to study due to their inherent variability, and the inconsistency of dedicated missions, but the longitudinal comparison between CRRES and VAP data, may enable such trends to be identified. Higher ambient solar wind driving might therefore have resulted in enhanced magnetotail loading and substorms injections. In addition, each of the above three points might also cause the plasmasphere/plasmaspheric hiss to extend to lower L shells during the CRRES era than during the VAP era. Lower energies are particularly affected by this loss mechanism, as found by the evolution of the energy spectra within VAP and CRRES observations [Zhao et al., 2019, Johnston et al., 2010].
@@
The bimodality in the decays rates is less apparent in the lifetimes beyond L* = 4, with the high Kp lifetimes now more often lower than the low Kp lifetimes and the average even lower than the VAP observations beyond L* = 5. The CRRES-MEA [Meredith et al., 2006] lifetimes are lower than the HEEF lifetimes in the outer belt between L = 2-4 although are higher than determined by VAP. The reason behind the differences with the MEA dataset is initially surprising given the HEEF and MEA datasets have been cross-calibrated [Johnston et al., 2014]. Comparisons between the Meredith et al. [2006] fits at the MEA energy channel of 0.604 MeV (not shown) and the HEEF energy channel of 0.65 MeV in Figure 5, indicate that the MEA algorithm is biased towards shorter fits and the HEEF fitting algorithms towards longer fits. This is evident near days 270, 495 and 520 in Figure 5e as these events feature some intrinsic variability in the fluxes, and Kp and DST indices, see Figure 1, but which overall are still well-approximated by periods of exponential decay if the algorithm is able to move past local minima in the Pearson values. This type of discrepancy may also explain further differences between the various studies reporting electron loss timescales.
@@
A further factor to be considered is that Brautigm and Bell [1995] adopt an energy span of 0.50-0.80 MeV for HEEF energy channel 0 and the electron lifetimes derived from HEEF for 0.65 MeV are therefore influenced by longer lifetime electrons than the MEA lifetimes. This is the energy-dependent decay seen in Meredith et al. [2006] and Zhao et al. [2019], i.e. as lower energy electrons are depleted the higher energy ones will dominate and may thus influence the lifetime more.
@@
It should be noted that Claudepierre et al. [2020] identify contamination due to Brehmstrahlung from higher energy electrons as a further potential cause of these longer lifetimes in the MEA and SAC/DEMETER datasets. As CRRES did not downlink the data required to identify and correct this effect, it is not possible to ascertain the extent to which HEEF was affected via the same method as VAP. However, we don’t tend to observe features at 0.65 MeV similar to those reported by Claudepierre et al. [2020] at 0.6 MeV in the VAP uncorrected data, which tend to show two different decay periods when the contamination is present, with a rapid initial decay and then a more gradual decay. This suggests that contamination from Bremsstrahlung is not so important for the HEEF data. The 0.65 MeV lifetimes also peak beyond L* = 4, whereas the multi-MeV lifetimes peak below L* = 3 and so the multi-MeV electrons are not directly correlated with the elevated lifetimes at 0.65 MeV. The HEEF and VAP lifetimes at higher energies also show good agreement, see subsequent discussion, and the HEEF lifetimes are actually shorter in the outer belt, consistent with higher accompanying geomagnetic activity.
@@
At 0.95 MeV, Figure 6b and 7b, the theory bounds many of the fitted lifetimes but is lower than the observations when scaling by geomagnetic activity with half the fitted lifetimes outside the upper bound predicted by theory. The HEEF, MEA and SAC/DEMETER observations however show similar lifetimes at all L shells. The VAP lifetimes appear slightly lower, consistent again with the lower geomagnetic activity argument discussed previously with reference to Figure 7a. At 1.60 MeV, Figure 6c and 7c, the HEEF and VAP observations show similar average values at L shells of 4.5-5.5 but further in, the HEEF results decay more slowly in-line with the previously discussed high Kp-associated decays. This trend is similar at 2.00 MeV, but the HEEF lifetimes are noticeably shorter than the VAP lifetimes beyond L shells of 4. The HEEF lifetimes are similar to the VAP lifetimes at 2.35, 3.75 and 3.15 MeV, but again do decay slightly faster at L shells greater than 4. The HEEF observations provide the first calculation of lifetimes at $>$4 MeV, and comparisons with previous observations at 4.55 MeV are therefore not possible. The fact that the VAP average lifetimes peak further out than the HEEF lifetimes may also in part be due to the VAP lifetimes using L instead of L* [e.g. Roederer and Lejosne, 2018].
@@
At L shells less than 4, discrepancies between theory and observations have been highlighted to potentially result from the lack of EMIC wave models [He et al., 2023] which preferentially affect larger energies, lower-frequency hiss [Ni et al., 2014] or wave-normal angle effects [Hartley et al., 2018]. The EMIC wave models used herein indeed only extend down to L* = 3.25, below which these waves are difficult to identify. The discrepancies between the observations and the theory could also well-derive from the assumption in Equation 1 of pure pitch angle diffusion. For example, Ross et al. [2021] employ 3-D modelling to better capture the decay of 2.5 MeV electrons near L = 3.5 using combined energy and radial diffusion. Su et al. [2012] indeed highlight that at L shells greater than 5, the loss timescales are largely insensitive to energy due to radial transport. It is therefore likely that radial diffusion also plays a role at lower L shells and particularly during higher levels of geomagnetic activity when the magnetopause may be more compressed.
Although the HEEF lifetimes are well-bounded by the theory, the lifetimes are not strictly ordered by geomagnetic activity, with data at lower L shells having been highlighted as corresponding to higher Kp indices. A spread in the lifetimes at higher L shells is less apparent but some of the intermediate decays associated with Kp values of 1-3 still not correctly ordering the decays. This suggests the varying role of alternative transport processes such as energy and radial diffusion.

### v2

@@ 5 HEEF Electron Lifetimes
To examine electron lifetimes throughout the CRRES era, Figure 5 then shows data across a range of L shells in the lowest energy channel, 0.65 MeV. The fits are once again overlaid on top of the data. The automated algorithm returns increasing numbers of fits at the intermediate L shells examined, overall visibly well-capturing exponentially decaying fluxes. A few flux decreases can however also be seen to be missed. In some instances, e.g. Figure 5d, it might appear that there are two stages to the decay process, for example in Figure 5g after day 450. These often appear as a shallow gradient followed by a greater decrease in the differential number flux. Further non-uniform decays are evident near the peak fluxes themselves over shorter time periods, with notably rounded structures at the lower L shells. This may in part be due to increasing production close to the initial event, varying geomagnetic activity levels during the decay, or potentially different eigenmodes dominating the pitch angle diffusion. At the lowest L shells, the electron fluxes decay over extended periods of tens-hundreds of days. The automated nature of the algorithm struggles with this extended period due to a shorter minimum fit window required to be defined to resolve the shorter timescales at larger L shells. The fluxes also exhibits significant short-term variability in the belt, on the order of several days. While the pitch angle spectra in Figure 2 appear coherent, we restrict our lifetime calculations in this region to a single event between day 265 and 310 at 0.65 MeV.
The decay periods are well-represented by the exponential fits in the region 3.0 $\leq$ L* $\leq$ 4.5. At L* $\geq$ 5, only a few fits are visible. This may in part be due to the algorithm being difficult to optimise for such a wide spread in lifetimes where, at these greater radial distances, further transport processes associated with radial diffusion [Su et al., 2010], drift shell splitting [Sibeck et al., 1987] and drift orbit bifurcations [Desai et al., 2021a] become significant.
Figure 6 displays the results across all L shells in the slot region and outer belt in each energy channel with each fit colored according to its mean Kp index across the entire decay period. This allows analysis of the conditions that exist during the quiet time following the event and determine the electron lifetimes. The average fit at each 0.1 L bin is also shown as a black line together with error bars representing the standard deviation. At the lowest L shells at the edge of the inner belt, the lifetimes of the 0.65 MeV electrons rapidly increase upward. The inner belt lifetimes are not discussed further here with regards to geomagnetic activity, due to their pitch angle dependence, but are discussed in the subsequent section. Decreased slot region lifetimes between L* = 2-3 appear at this energy, as in Figure 3, before the lifetimes increase to another smaller maximum near L = 4.
@@
The lifetimes of the 0.95 MeV electrons are lower than those of the 0.65 MeV electrons near L* = 4.0 but this trend does not extend into the slot region where the reverse is true. There is a further peak near L = 2.5-3 but this is due to a large spread in the data here. At energies greater than 1 MeV, the peak lifetime in the outer belt moves inward with increasing energy from L* $\approx$ 4.5 at 1 MeV down to L* $\approx$ 3 at 4.55 MeV. The location of the shortest lifetimes in the slot region fluxes also move inwards, with increasing energies displaying longer lifetimes.
@@
In addition to the average trends, the Kp values of each fitted lifetime indicate how the decays scale with geomagnetic activity. At higher L shells, i.e. greater than L* = 3, elevated geomagnetic activity appears to drive fast losses with some of the shortest timescales showing Kp values near 5. At the lower L’s, however, and particularly at the lowest energy channel of 0.65 MeV, high Kp values often correspond to longer decay timescales. High Kp scenarios will still result in injections from the tail and subsequent acceleration of energetic particles and this trend therefore might be explained by sustained production of electrons occurring alongside pitch angle scattering, a phenomenon which preferentially affects lower energies. This is further discussed in Section 5.
@@
Figure 7: Experimental and theoretical electron lifetimes as a function of L shell in the outer belt and slot region for (a-h) 0.65, 0.95, 1.60, 2.00, 2.35, 2.75, 3.15 and 3.75 MeV electrons respectively. The black, red, grey, green and pink traces represent the experimental lifetimes derived from the CRRES-HEEF (this study), CRRES-MEA also from cycle 22 [Meredith et al., 2006], VAP from cycle 24 [Claudepierre et al., 2020], and SAC/DEMETER [Benck et al., 2010] and SAMPEX [Meredith et al., 2009] datasets from cycle 23. The dashed blue and orange traces represent the pitch angle diffusion theoretical lifetimes for quiet and high geomagnetic activity, respectively. All data is interpolated to the HEEF energies, whereas the SAMPEX 2–6 MeV data are associated with the 2.35 MeV HEEF channel.
@@ 6 Comparisons & Discussion
In Figure 6, the pitch angle diffusion theory matches the observed trend of increased lifetimes near the inner belt at 0.65 MeV and depressed lifetimes in the slot region. In the inner part of the outer radiation belt, the quiet-time theoretical lifetimes at this energy are a factor of two or more smaller than the observed lifetimes in the region 3.2 <L<<L*< 4.6. These larger HEEF lifetimes are similar to those observed by SAC/DEMETER in cycle 23 [Benck et al., 2010] and also by ATS-1 [Vampola, 1971] in cycle 20 see Claudepierre et al. [2020, Fig. 2f therein]. This however is not observed in the VAP lifetimes. Further examination of Figure 6 reveals a cause of this discrepancy. The underlying HEEF fits, most visibly between L* = 3-4, feature a clustering of lifetimes at the lower lifetimes ranges, similar to those reported by the MEA and closer to those observed by VAP. However, an additional clustering above this demonstrates a bi-modality within the dataset, with this second cluster corresponding to higher Kp values. This bimodality is also identifiable in the underlying fits on Figure 5e,f where many of the fitted lifetimes have shallower gradients.
@@
The consequent differences in the average lifetimes in Figure 7a may therefore have been caused by differing levels of geomagnetic activity, with CRRES having observed an average Kp index of 2.67 compared to the VAP average of 1.67. This can be explained through three distinct mechanisms. Firstly, the intensity of geomagnetic storms were significantly greater in solar cycle 22 compared with solar cycle 24. These, during the latter half of the mission, indeed appear to be associated with the slower decays in Figure 5. These intense and longer lasting events influence the average Kp indices of a decay and thus skew the decay rates to higher Kp values, producing a “left-right" Kp bias as indeed shown by Meredith et al. [2006, Figure 6 therein]. Secondly, disturbed geomagnetic conditions would more frequently reoccur during a decay period. This is visibly consistent with several of the decays observed, see for example the decays days 270 and 295 in Figure 5f-g where large injections/enhancements occur during the decay period. Thirdly, it is well-known that typical solar wind parameters have declined significantly over the past 50 years, with differences between the 1970-1990’s and 2010’s having been highlighted as high as tens to over fifty percent [McComas et al., 2013]. Such long-term trends in the radiation belts are difficult to study due to their inherent variability, and the inconsistency of dedicated missions, but the longitudinal comparison between CRRES and VAP data, may enable such trends to be identified. Higher ambient solar wind driving might therefore have resulted in enhanced background levels of magnetotail loading and substorms injections. In addition, each of the above three points might also cause the plasmasphere/plasmaspheric hiss to extend to lower L shells during the CRRES era than during the VAP era. Lower energies are particularly affected by this loss mechanism, as found by the evolution of the energy spectra within VAP and CRRES observations [Zhao et al., 2019, Johnston et al., 2010].
@@
The bimodality in the decays rates is less apparent in the lifetimes beyond L* = 4, with the high Kp lifetimes now more often lower than the low Kp lifetimes and the average even lower than the VAP observations beyond L* = 5. The CRRES-MEA [Meredith et al., 2006] lifetimes are lower than the HEEF lifetimes in the outer belt between L = 2-4 although are higher than determined by VAP. The reason behind the differences with the MEA dataset is initially surprising given the HEEF and MEA datasets have been cross-calibrated [Johnston et al., 2014b]. Comparisons between the Meredith et al. [2006] fits at the MEA energy channel of 0.604 MeV (not shown) and the HEEF energy channel of 0.65 MeV in Figure 5, indicate that the HEEF fitting algorithm skews towards longer fits. This is evident near days 270, 495 and 520 in Figure 5f as these events feature some intrinsic variability in the fluxes, and Kp and DST indices, see Figure 1, but which overall are still well-approximated by periods of exponential decay if the algorithm is able to move past local minima in the Pearson values. This type of discrepancy may also explain further differences between the various studies reporting electron loss timescales.
@@
A further factor to be considered is that Brautigam and Bell [1995] adopt an energy span of 0.50-0.80 MeV for HEEF energy channel 0 and the electron lifetimes derived from HEEF for 0.65 MeV are therefore influenced by longer lifetime electrons than the MEA lifetimes. This is the energy-dependent decay seen in Meredith et al. [2006] and Zhao et al. [2019], i.e. as lower energy electrons are depleted the higher energy ones will dominate and may thus influence the lifetime more.
@@
It should be noted that Claudepierre et al. [2020] identify contamination due to Brehmstrahlung from higher energy electrons as a further potential cause of these longer lifetimes in the MEA and SAC/DEMETER datasets. As CRRES did not downlink the data required to identify and correct this effect, it is not possible to ascertain the extent to which HEEF was affected via the same method as VAP. However, we don’t tend to observe features at 0.65 MeV similar to those reported by Claudepierre et al. [2020] at 0.6 MeV in the VAP uncorrected data, which tend to show two different decay periods when the contamination is present, with a rapid initial decay and then a more gradual decay. This suggests that contamination from Bremsstrahlung is not so significant for the HEEF data. The 0.65 MeV lifetimes also peak beyond L* = 4, whereas the multi-MeV lifetimes peak below L* = 3 and so the multi-MeV electrons are not directly correlated with the elevated lifetimes at 0.65 MeV. The HEEF and VAP lifetimes at higher energies also show good agreement, see subsequent discussion, and the HEEF lifetimes are actually shorter in the outer belt, consistent with higher accompanying geomagnetic activity.
In the inner belt, the decay timescales match closely with the VAP decay timescales, which were also determined for 90 degrees flux. These elevated fluxes in the HEEF dataset, see Figure 5a, appear to have been caused by storms at or prior to the start of the mission followed by relatively quiet conditions which allow their decay timescales to be calculated across this period. These loss timescales approach 200 days and is also evident in that during the second half of the mission, see again Figure 5a where geomagnetic activity was higher, the 0.65 MeV flux is unable to decay and appears to be largely on an upward trend from around day 450 to 550. This points to the controlling influence of slot region fluxes on electron flux levels in the inner belt.
@@
At 0.95 MeV, Figure 6b and 7b, the theory bounds many of the fitted lifetimes but is lower than the observations when scaling by geomagnetic activity with half the fitted lifetimes outside the upper bound predicted by theory. The HEEF, MEA and SAC/DEMETER observations however show similar lifetimes at all L shells. The VAP lifetimes appear slightly lower, consistent again with the lower geomagnetic activity argument discussed previously with reference to Figure 7a. At 1.60 MeV, Figure 6c and 7c, the HEEF and VAP observations show similar average values at L shells of 4.5-5.5 but further in, the HEEF results decay more slowly in-line with the previously discussed high Kp-associated decays. This trend is similar at 2.00 MeV, but the HEEF lifetimes are noticeably shorter than the VAP lifetimes beyond L shells of 4. The HEEF lifetimes are similar to the VAP lifetimes at 2.35, 3.75 and 3.15 MeV, but again do decay slightly faster at L shells greater than 4. The HEEF observations provide the first calculation of lifetimes at >>4 MeV, and comparisons with previous observations at 4.55 MeV are therefore not possible. The fact that the VAP average lifetimes peak further out than the HEEF lifetimes may also in part be due to the VAP lifetimes using L instead of L* [e.g. Roederer and Lejosne, 2018].
@@
At L shells less than 4, discrepancies between theory and observations have been highlighted to potentially result from the lack of EMIC wave models [He et al., 2023] which preferentially affect larger energies, lower-frequency hiss [Ni et al., 2014] or wave-normal angle effects [Hartley et al., 2018]. The EMIC wave models used herein indeed only extend down to L* = 3.25, below which these waves are difficult to identify. The discrepancies between the observations and the theory could also well-derive from the assumption in Equation 1 of pure pitch angle diffusion. For example, Ross et al. [2021] employ 3-D modelling to better capture the decay of 2.5 MeV electrons near L = 3.5 using combined energy and radial diffusion. Su et al. [2012] indeed highlight that at L shells greater than 5, the loss timescales are largely insensitive to energy due to radial transport. It is therefore likely that radial diffusion also plays a role at lower L shells and particularly during higher levels of geomagnetic activity when the magnetopause may be more compressed. Broll et al. [2023] also shows that simulated pitch angle diffusion evolves differently than empirical estimates when non-equilibrium pitch angle distributions are considered, and indeed shows this approach yield decay timescales much closer to the observed timescales than the theoretical ‘lifetimes’ (from $D_{\alpha\alpha}$’s slowest-decaying mode) in the inner belt.
Although the HEEF lifetimes are well-bounded by the theory, the lifetimes are not strictly ordered by geomagnetic activity, with decays at lower L shells having been highlighted as corresponding to higher Kp indices. An event specific spread in the lifetimes at higher L shells is less apparent but some of the intermediate decays associated with Kp values of 1-3 are still not correctly ordering the decays. This suggests the event specific nature of the decays [Ripoll et al., 2016], potential role of alternative transport processes such as energy and radial diffusion, and errors with using a single global geomagnetic activity index for parametrising radiation belt dynamics.